OpenStudy (anonymous):

Solve and describe the steps you took to solve: 5(n-5)>2(n+6)

OpenStudy (anonymous):

you're going to vvant to solve for each side first. multiply vvhats outside the parenthesis and inside. you'll get 5n - 25 >2n + 12

OpenStudy (anonymous):

you have to use distributive property. 5(n-5) = 5n-25 because you multiply 5 by n and -5 2(n+6) = 2n+12 because you multiply 2 by n and 6 then you have 5n-25>2n+12 then combine like terms: 5n-2n= 3n 12+25=37 3n>37 divide 3 by both sides and you should end up with n>12.3333.... forgive me if im wrong but im almost positive

OpenStudy (anonymous):

Ah, thank you! I see!

OpenStudy (anonymous):

youre welcome!(: